# dyGRASS: Dynamic Spectral Graph Sparsification via Localized Random Walks on GPUs

### v1

@@ Abstract
This work presents dyGRASS, an efficient dynamic algorithm for spectral sparsification of large undirected graphs that undergo streaming edge insertions and deletions. At its core, dyGRASS employs a random-walk-based method to efficiently estimate node-to-node distances in both the original graph (for decremental update) and its sparsifier (for incremental update). For incremental updates, dyGRASS enables the identification of spectrally critical edges among the updates to capture the latest structural changes. For decremental updates, dyGRASS facilitates the recovery of important edges from the original graph back into the sparsifier. To further enhance computational efficiency, dyGRASS employs a GPU-based non-backtracking random walk scheme that allows multiple walkers to operate simultaneously across various target updates. This parallelization significantly improves both the performance and scalability of the proposed dyGRASS framework. Our comprehensive experimental evaluations reveal that dyGRASS achieves approximately a 10$\times$ speedup compared to the state-of-the-art incremental sparsification (inGRASS) algorithm while eliminating the setup overhead and improving solution quality in incremental spectral sparsification tasks. Moreover, dyGRASS delivers high efficiency and superior solution quality for fully dynamic graph sparsification, accommodating both edge insertions and deletions across a diverse range of graph instances originating from integrated circuit simulations, finite element analysis, and social networks.
@@ I Introduction
    By accelerating random walks on Graphics Processing Units (GPUs), dyGRASS achieves a 10×10$\times$10 × runtime speedup over the CPU-based implementation and 2×2$\times$2 × speedup over the state-of-the-art method [32] on real-world large graphs, while preserving solution quality.
@@ II-A Graph Laplacian, quadratic form, and spectral similarity
For a weighted undirected graph $G=(V,E,w)$, $V$ ($|V|=n$) and $E$ represent the sets of vertices and edges, respectively, and $w$ denotes a positive weight function with $w_{i,j}$ denoting the edge weight. Let $A$ denote the $n\times n$ adjacency matrix, then the Laplacian matrix $L_{G}$ is defined as $L_{G}:=D-A$, where $D$ is the degree matrix. The eigenvalues and eigenvectors of $L_{G}$ offer insights into graph structure properties such as the number of connected components and vertex connectivity. The Laplacian quadratic form $x^{\top}L_{G}x$ is a fundamental tool in spectral graph theory, enabling analysis of graph cuts, clustering, and conductance.
Spectral similarity between the Laplacians of graphs $A$ and $B$ can be quantified using the following inequality [33]:
$\frac{x^{\top}L_{A}x}{\epsilon}\leq x^{\top}L_{B}x\leq\epsilon x^{\top}L_{A}x,$ (1)
where $L_{A}$ and $L_{B}$ represent the Laplacian matrices of $A$ and $B$, with a parameter $\epsilon>0$. A smaller $\epsilon$ or relative condition number $\kappa(L_{A},L_{B})$ indicates greater spectral similarity between $A$ and $B$.
@@ II-C Effective resistance
$R^{eff}_{p,q}=b_{p,q}^{\top}L^{\dagger}b_{p,q}=\sum\limits_{i=2}^{|V|}\frac{(u_{i}^{\top}b_{p,q})^{2}}{u_{i}^{\top}Lu_{i}},$ (2)
where $L^{\dagger}$ denotes the Moore-Penrose pseudo-inverse of the graph Laplacian matrix $L$, $u_{i}\in\mathbb{R}^{|V|}$ for $i=1,...,|V|$ denote the unit-length, mutually-orthogonal eigenvectors corresponding to Laplacian eigenvalues $\lambda_{i}$ for $i=1,...,|V|$, ${b_{p}}\in\mathbb{R}^{|V|}$ denotes the standard basis vector with all zero entries, with the $p$-th entry equal to 1, and ${b_{p,q}}=b_{p}-b_{q}$.
@@ II-D Random walk
A random walk on a graph is a stochastic process where a walker starts at a vertex and moves to a randomly chosen neighbor at each step. Random walks are a fundamental tool in graph analysis, offering insights into structural and spectral properties and enabling sampling, distance estimation, and connectivity analysis without full graph exploration [34]. Let $G=(V,E)$ be an undirected graph, with $|V|=n$ vertices and $|E|=m$ edges. For any two vertices $p,q\in V$, the hitting time $H_{p,q}$ is the expected number of steps for a walker starting at $p$ to reach $q$ for the first time.
@@
For any two vertices $p$ and $q$ in $G$, the commute time $C_{p,q}$ is the expected round-trip time between $p$ and $q$, which is directly proportional to the effective resistance distance between $p$ and $q$: $C_{p,q}=H_{p,q}+H_{q,p}=2m\cdot R^{eff}_{p,q}$ [35].
@@ III-B Incremental spectral sparsification with dyGRASS
(a) Spectral Impact of Edge Insertions: Suppose an edge $e_{p,q}\notin E^{(l)}$ is newly introduced between vertices $p$, $q$ during the $l$-th iteration, updating the original graph $G^{(l)}=(V,E^{(l)},w^{(l)})$. When incorporating $e_{p,q}$ into current sparsifier $H^{(l)}=(V,E^{\prime(l)},w^{\prime(l)})$, the eigenvalue perturbation for each eigenvector of $H^{(l)}$ can be assessed by the following lemma [26].
@@
Let $H=(V,E^{\prime},w^{\prime})$, where $w^{\prime}:E^{\prime}\rightarrow\mathbb{R}+$, denote the sparsified weighted graph of $G$, and $L_{H}$ denotes its Laplacian matrix. The $i$-th Laplacian eigenvalue perturbation due to $\delta L_{H}=w^{\prime}_{p,q}b_{pq}b^{\top}_{pq}$ can be computed as:
@@
$\Delta_{K}=\sum\limits_{i=2}^{{K}}\frac{\delta{{\lambda}_{i}}}{\lambda_{i}}=w_{p,q}\|U_{K}^{\top}b_{pq}\|^{2}_{2}\approx w_{p,q}R(p,q).$ (5)
Lemma 3.2 defines the spectral distortion $\Delta_{K}$ of an edge as the product of its edge weight and effective resistance between its two end vertices. We next show that the overall distortion remains bounded by the condition number $\kappa(L_{G},L_{H})$.
Consider the generalized eigenvalue problem for the matrix pencil $(L_{H},L_{G})$ and apply the Courant-Fischer theorem. Specifically, we have:
@@
where $b_{p,q}$ is the incidence vector of edge $e_{p,q}$, and $\mathbf{1}$ is the all-one vector. From
$\frac{b_{p,q}^{\top}L_{H}^{+}\,b_{p,q}}{b_{p,q}^{\top}L_{G}^{+}\,b_{p,q}}\;=\;\frac{R_{p,q,H}}{R_{p,q,G}},$ (7)
we see that taking the maximum over all node pairs $(p,q)$ implies
$\max_{p,q}\,\frac{R_{p,q,H}}{R_{p,q,G}}\;\;\leq\;\;\lambda_{\max}(L_{G}\,L_{H}^{+}).$ (8)
@@
$\omega_{p,q}\,R_{p,q,H}\;\leq\;\frac{R_{p,q,H}}{R_{p,q,G}}\;\leq\;\kappa(L_{G},L_{H}),\quad\forall\,p,q\in V.$ (11)
@@
Equation (5) indicates that an edge $e_{p,q}$ with a large product $w_{p,q}\,R_{p,q}$ can impose significant spectral distortion $\Delta_{K}$. Yet, directly computing the effective resistance $R_{p,q}$ for every inserted edge is prohibitively expensive on large graphs. To circumvent this bottleneck, we adopt a non-backtracking random walk approach that efficiently approximates or upper-bounds $R_{p,q}$ to enable rapid edge filtering.
(b) Non-Backtracking Random Walks and Approximate Resistance Distance: At the core of our method lies the observation that the accumulated effective resistance along the shortest path between two vertices $p$ and $q$ provides an upper bound on the true effective resistance $R_{p,q}$. Specifically, if one could identify the shortest path $P^{*}$ from $p$ to $q$, then the sum $\sum_{e\in P^{*}}r_{e}$ (where $r_{e}=1/w_{e}$) would serve as a valid upper bound on $R_{p,q}$, as resistance in parallel can only reduce total resistance.
However, directly computing shortest paths or all-pairs effective resistance remains computationally prohibitive in large-scale graphs. To circumvent this, we adopt a non-backtracking random walk (NBRW) approach, which probabilistically explores paths from $p$ without immediately revisiting the previous node, emulating a depth-first search pattern but in a randomized and lightweight manner. On tree-structured graphs, the path generated by NBRW from $p$ to $q$ corresponds exactly to the unique shortest path, making the accumulated resistance along the walk equal to $R_{p,q}$.
@@
This Monte Carlo strategy enables a practical, localized approximation of the effective resistance. The approximate spectral distortion for edge $e_{p,q}$ is then defined as $\tilde{\Delta}=\omega_{p,q}\tilde{R}_{p,q}$, and a distortion threshold $K$ is imposed to guide the random walk: walkers terminate their search once the cumulative approximate distortion exceeds this threshold ($\omega_{p,q}\tilde{R}_{\text{cumulative}}>K$), dynamically adjusting their search range rather than using a fixed step limit. In other words, the threshold $K$ acts as an effective proxy to filter edges in the sparsification process. For a newly introduced edge $e_{p,q}$:
  • If vertex $q$ is not reachable from vertex $p$ within the approximate distortion threshold $K$, the edge $e_{p,q}$ is considered spectrally critical and should be included in the sparsifier.
  • If vertex $q$ is reachable within the approximate distortion threshold $K$, the edge $e_{p,q}$ can be safely pruned without significantly impacting spectral similarity.
@@ III-D The dyGRASS algorithm flow
1:  Input: $G^{(0)}$: Original graph, $H^{(0)}$: Sparsifier, EdgeStream: Stream of updates, distortion threshold $K$, maximum step size $T$, number of walkers $s$, update type $choice$
@@ IV-A Dynamic spectral sparsification using dyGRASS
We evaluate the performance of dyGRASS by comparing the off-tree edge density of the sparsifier, which is defined as $D_{H}:=\frac{|E|}{|V|}-1$ the same relative condition number is achieved using inGRASS and dyGRASS methods. The relative condition number $\kappa(L_{G},L_{H})$ can be deployed to quantify the spectral similarity between the original graph $G$ and its graph sparsifier $H$ [11, 12]. A smaller relative condition number implies higher spectral similarity between the graph $G$ and its sparsifier $H$.
@@ IV-B Sparsifier densities after multiple incremental updates
Fig. 5 illustrates dyGRASS’s dynamic updating capability in ten iterations for handling the circuit graph ”G2_circuit”. Since dyGRASS can capture the latest structural changes (e.g., decreasing resistance diameters) in every update iteration via localized random walks, it allows the newly inserted edges to be selected more optimally considering the most recent updates. As a result, dyGRASS is able to achieve much lower sparsifier density than inGRASS as more updates are made. In contrast, when more edges are inserted into the graph, the resistance distances between nodes will monotonically decrease, leading to an increasing overestimation of the spectral distortion of each new edge since inGRASS keeps processing newly inserted edges using the dated resistance embedding.
@@
$\kappa(L_{G}^{(0)},L_{H}^{(0)})$ $D_{H}^{(0)}$Iters $\Delta D_{G}^{(10)}$Iters $D_{H}^{(10)}$Iters
@@ IV-C Preconditioned Solver Iterations
To solve a linear system $L_{G}\,x=b$ more efficiently, we use a spectral sparsifier $H$ as a preconditioner. That is, we replace $L_{G}$ by $L_{H}$ in each iterative step, effectively solving $L_{H}^{-1}L_{G}\,x=L_{H}^{-1}b$. The convergence rate hinges on the condition number $\kappa(L_{G},L_{H})$, so a high-quality sparsifier leads to fewer solver iterations. In this experiemnt, the PCG solver terminates when the relative residual norm $<10^{-8}$.
@@ IV-D Runtime scalability of dyGRASS
In Fig. 6, we compare the runtime scalability of dyGRASS and inGRASS across various real-world graphs. Over ten iterative incremental updates, the GPU implementation of dyGRASS achieves approximately a 70$\times$ speedup over its CPU counterpart. Compared to inGRASS, dyGRASS offers about a 10$\times$ improvement for large graphs. Although dyGRASS maintains $O(1)$ time complexity per edge update, the total runtime increases with graph size due to the growing number of updates. Specifically, we fix the update density at approximately 25% of the number of vertices. As a result, larger graphs involve proportionally more edge updates per iteration, leading to an increase in total update time.

### v3

@@ Abstract
[…]
@@ I Introduction
    By accelerating random walks on Graphics Processing Units (GPUs), dyGRASS achieves a $70\times$ runtime speedup over the CPU-based implementation and at most $10\times$ speedup over the state-of-the-art method [32] on real-world large graphs, while preserving solution quality.
@@ II-A Graph Laplacian, quadratic form, and spectral similarity
For a weighted undirected graph $G=(V,E,w)$, $V$ ($|V|=n$) and $E$ ($|E|=m$) represent the sets of vertices and edges, respectively, and $w$ denotes a positive weight function with $w_{i,j}$ denoting the edge weight. Let $A$ denote the $n\times n$ adjacency matrix, then the Laplacian matrix $L$ is defined as L:=DAL:=D-A, where $D$ is the degree matrix. The eigenvalues and eigenvectors of $L$ offer insights into graph structure properties such as the number of connected components and vertex connectivity. The Laplacian quadratic form $x^{\top}Lx$ is a fundamental tool in spectral graph theory, enabling analysis of graph cuts, clustering, and conductance.
Spectral similarity between the Laplacians of graphs $G$ and $H$ can be quantified using the following inequality [33]:
$\frac{x^{\top}L_{G}x}{\epsilon}\leq x^{\top}L_{H}x\leq\epsilon x^{\top}L_{G}x,$ (1)
where $L_{G}$ and $L_{H}$ represent the Laplacian matrices of $G$ and $H$, with a parameter $\epsilon>0$. A smaller $\epsilon$ or relative condition number $\kappa(L_{G},L_{H})$ indicates greater spectral similarity between $G$ and $H$.
@@ II-C Effective resistance
$R^{\text{eff}}_{p,q}:=b_{p,q}^{\top}L^{\dagger}b_{p,q}=\sum\limits_{i=2}^{|V|}\frac{(u_{i}^{\top}b_{p,q})^{2}}{u_{i}^{\top}Lu_{i}},$ (2)
where $L^{\dagger}$ denotes the Moore-Penrose pseudo-inverse of the graph Laplacian matrix $L$, $u_{i}\in\mathbb{R}^{|V|}$ for i=1,,|V|i=1,...,|V| denote the unit-length, mutually-orthogonal eigenvectors corresponding to Laplacian eigenvalues $\lambda_{i}$ for i=1,,|V|i=1,...,|V|, ${b_{p}}\in\mathbb{R}^{|V|}$ denotes the standard basis vector with all zero entries, with the $p$-th entry equal to 1, and ${b_{p,q}}=b_{p}-b_{q}$.
@@ II-D Random walk
A random walk on a graph is a stochastic process where a walker starts at a vertex and moves to a randomly chosen neighbor at each step. Random walks are a fundamental tool in graph analysis, offering insights into structural and spectral properties and enabling sampling, distance estimation, and connectivity analysis without full graph exploration [34]. For any two vertices $p,q\in V$, the hitting time $h_{p,q}$ is the expected number of steps for a walker starting at $p$ to reach $q$ for the first time.
@@
For any two vertices $p$ and $q$ in $G$, the commute time $C_{p,q}$ is the expected round-trip time between $p$ and $q$, which is directly proportional to the effective resistance distance between $p$ and $q$: $C_{p,q}=h_{p,q}+h_{q,p}=2m\cdot R^{\text{eff}}_{p,q}$ [35].
@@ III-B Incremental spectral sparsification with dyGRASS
Spectral Impact of Edge Insertions: Suppose an edge $e_{p,q}\notin E^{(l)}$ is newly introduced between vertices $p$, $q$ during the $l$-th iteration, updating the original graph $G^{(l)}=(V,E^{(l)},w^{(l)})$. When incorporating $e_{p,q}$ into current sparsifier $H^{(l)}=(V,E^{\prime(l)},w^{\prime(l)})$, the eigenvalue perturbation for each eigenvector of $H^{(l)}$ can be assessed by the following lemma [26].
@@
Let $H=(V,E^{\prime},w^{\prime})$, where $w^{\prime}:E^{\prime}\rightarrow\mathbb{R}+$, denote the sparsified weighted graph of $G$, and $L_{H}$ denotes its Laplacian matrix. The $i$-th Laplacian eigenvalue perturbation $\delta\lambda_{i}$ due to $\delta L_{H}=w^{\prime}_{p,q}b_{pq}b^{\top}_{pq}$ can be computed as:
@@
$\Delta_{K}:=\sum\limits_{i=2}^{{K}}\frac{\delta{{\lambda}_{i}}}{\lambda_{i}}=w_{p,q}\|U_{K}^{\top}b_{pq}\|^{2}_{2}\approx w_{p,q}R^{\text{eff}}_{p,q}.$ (5)
Lemma 3.2 defines the spectral distortion $\Delta_{K}$ of an edge as the product of its edge weight and effective resistance between its two end vertices. We next show that the all distortion remains bounded by the condition number $\kappa(L_{G},L_{H})$.
For a graph $G$ and a subgraph $H$ (obtained by deleting edges), consider the generalized eigenvalue problem for the matrix pencil $(L_{H},L_{G})$ and apply the Courant-Fischer theorem. Specifically, we have:
@@
where $\mathbf{1}$ is the all-one vector.
$\frac{b_{p,q}^{\top}L_{H}^{+}\,b_{p,q}}{b_{p,q}^{\top}L_{G}^{+}\,b_{p,q}}\;=\;\frac{R^{\text{eff}}_{p,q,H}}{R^{\text{eff}}_{p,q,G}},$ (7)
From equations (6) and (7), we see that taking the maximum over all node pairs $(p,q)$ implies
$\max_{p,q}\,\frac{R^{\text{eff}}_{p,q,H}}{R^{\text{eff}}_{p,q,G}}\;\;\leq\;\;\lambda_{\max}(L_{G}\,L_{H}^{+}).$ (8)
@@
$\omega_{p,q}\,R^{\text{eff}}_{p,q,H}\;\leq\;\frac{R^{\text{eff}}_{p,q,H}}{R^{\text{eff}}_{p,q,G}}\;\leq\;\kappa(L_{G},L_{H}),\quad\forall\,p,q\in V.$ (11)
@@
Equation (5) indicates that an edge $e_{p,q}$ with a large product $w_{p,q}\,R^{\text{eff}}_{p,q}$ can impose significant spectral distortion $\Delta_{K}$. Yet, directly computing the $R^{\text{eff}}_{p,q}$ for every inserted edge is prohibitively expensive on large graphs. To circumvent this bottleneck, we adopt a non-backtracking random walk approach that efficiently approximates or upper-bounds $R^{\text{eff}}_{p,q}$ to enable rapid edge filtering.
Non-Backtracking Random Walks and Approximate Resistance Distance: At the core of our method lies the observation that the accumulated effective resistance along the shortest path between two vertices $p$ and $q$ provides an upper bound on the true $R^{\text{eff}}_{p,q}$. Specifically, if one could identify the shortest path $P^{*}$ from $p$ to $q$, then the sum $\sum_{e\in P^{*}}r_{e}$ (where $r_{e}=1/w_{e}$) would serve as a valid upper bound on $R^{\text{eff}}_{p,q}$, as resistance in parallel can only reduce total resistance.
However, directly computing shortest paths or all-pairs effective resistance remains computationally prohibitive in large-scale graphs. To circumvent this, we adopt a non-backtracking random walk (NBRW) approach, which probabilistically explores paths from $p$ without immediately revisiting the previous node, emulating a depth-first search pattern but in a randomized and lightweight manner. On a tree-structured graph, the path generated by NBRW from $p$ to $q$ corresponds exactly to the unique shortest path, making the accumulated resistance along the walk equal to $R^{\text{eff}}_{p,q}$.
@@
This Monte Carlo strategy enables a practical, localized approximation of the effective resistance. The approximate distortion is then defined as $\tilde{\Delta}:=\omega_{p,q}\tilde{R}_{p,q}$, and a distortion threshold $K$ is imposed to guide the random walk: walkers terminate their search once the cumulative approximate distortion exceeds this threshold ($\omega_{p,q}\tilde{R}_{\text{cumulative}}>K$), dynamically adjusting their search range rather than using a fixed step limit. In other words, the threshold $K$ acts as an effective proxy to filter edges in the sparsification process. For a newly introduced edge $e_{p,q}$:
  • If vertex $q$ is not reachable from vertex $p$ within the distortion threshold $K$, the edge $e_{p,q}$ is considered spectrally critical and should be included in the sparsifier.
  • If vertex $q$ is reachable within the distortion threshold $K$, the edge $e_{p,q}$ can be safely pruned without significantly impacting spectral similarity.
@@ III-D The dyGRASS algorithm flow
1:  Input: $G^{(0)}$: Original graph, $H^{(0)}$: Sparsifier, EdgeStream: Stream of updates, distortion threshold $K$, maximum step size $T$, number of walkers $s$, update type choicechoice
@@ IV-A Dynamic spectral sparsification using dyGRASS
Consistent with the prior inGRASS [32] study, we reuse its 14 benchmark graphs and replay the identical update stream: ten insertion batches that together add $\approx\!25\%$ of $|V|$ edges. We evaluate the performance of dyGRASS by comparing the off-tree edge density of the sparsifier, which is defined as $d_{H}:=\frac{|E|}{|V|}-1$ the same relative condition number is achieved using inGRASS and dyGRASS methods. The relative condition number $\kappa(L_{G},L_{H})$ can be deployed to quantify the spectral similarity between the original graph $G$ and its graph sparsifier $H$ [11, 12]. A smaller relative condition number implies higher spectral similarity between the graph $G$ and its sparsifier $H$.
@@ IV-B Sparsifier densities after multiple incremental updates
Fig. 5 illustrates dyGRASS’s dynamic updating capability in ten iterations for handling the circuit graph “G2_circuit”. Since dyGRASS can capture the latest structural changes (e.g., decreasing resistance diameters) in every update iteration via localized random walks, it allows the newly inserted edges to be selected more optimally considering the most recent updates. As a result, dyGRASS is able to achieve much lower sparsifier density than inGRASS as more updates are made. In contrast, when more edges are inserted into the graph, the resistance distances between nodes will monotonically decrease, leading to an increasing overestimation of the spectral distortion of each new edge since inGRASS keeps processing newly inserted edges using the dated resistance embedding.
@@
$\kappa(L_{G}^{(0)},L_{H}^{(0)})$ $d_{H}^{(0)}$ #Iters $\Delta d_{G}^{(10)}$ #Iters $d_{H}^{(10)}$ #Iters
@@ IV-C Preconditioned Solver Iterations
To solve a linear system $L_{G}\,x=b$ more efficiently, we use a spectral sparsifier $H$ as a preconditioner. That is, we replace $L_{G}$ by $L_{H}$ in each iterative step, effectively solving $L_{H}^{-1}L_{G}\,x=L_{H}^{-1}b$. The convergence rate hinges on the condition number $\kappa(L_{G},L_{H})$, so a high-quality sparsifier leads to fewer solver iterations. In this experiemnt, the PCG solver terminates when the relative residual norm $10^{-8}$.
@@ IV-D Runtime scalability of dyGRASS
[…]
@@ V Conclusion
VI Acknowledgments
This work is supported in part by the National Science Foundation under Grants CCF-2417619 and CCF-2212370.